# Huan Zhang

### v1

@@ Abstract
This paper develops a primal-dual dynamical system where the coefficients are designed in closed-loop way for solving a convex optimization problem with linear equality constraints. We first introduce a “second-order primal” + “first-order dual” continuous-time dynamical system, in which both the time scaling and Hessian-driven damping are governed by a feedback control of the gradient for the Lagrangian function. This system achieves the fast convergence rates for the primal-dual gap, the feasibility violation, and the objective residual along its trajectory. Subsequently, by time discretization of this system, we develop an accelerated primal-dual algorithm with a gradient-defined adaptive step size. We also obtain convergence rates for the primal-dual gap, the feasibility violation, and the objective residual. Furthermore, we provide numerical results to demonstrate the practical efficacy and superior performance of the proposed algorithm.
@@ 1.2 Motivation and related works
Over the past decades, a series of studies have explored second-order dynamical systems in continuous and discrete time for the unconstrained convex optimization problem
@@
where $\alpha\geq 3$, and obtained the $\mathcal{O}\left(\frac{1}{t^{2}}\right)$ convergence rate of the objective residual along the trajectory generated by the system (6). The system (6) has further been studied in several papers, including may2017 ; attouch2018 ; attouch2019 . They showed that the objective residual converges at a rate of order $o\left(\frac{1}{t^{2}}\right)$ for $\alpha>3$, and the trajectory weakly converges to the global minimizer of the objective function. In order to improve the convergence rate, Attouch et al. a2019 proposed the system (6) with the time scaling:
@@
where $\alpha\geq 1$ and $\beta:[t_{0},+\infty)\rightarrow(0,+\infty)$ is the time scaling function which is non-decreasing and continuously differentiable. They obtained the $\mathcal{O}\left(\frac{1}{t^{2}\beta(t)}\right)$ convergence rate of the objective residual along the trajectory generated by this system, and proposed an inertial proximal algorithm by time discretization of this system.
On the other hand, primal-dual dynamics methods were also attracted an increasing interest by many researchers for solving the linear equality constrained convex optimization problem (3). Over the past few years, there have been a wide variety of works devoted to “second-order primal” + “second-order dual” continuous-time dynamical systems for solving Problem (3), see bot2021n ; hehu2021 ; zeng2023 ; hn2023 . By time discretization of continuous-time dynamical systems, new accelerated primal-dual algorithms for solving Problem (3) have been proposed in hehu2022 ; bot2023 ; zhu2024 ; dingliu2025 . Since the computational cost of a primal-dual method primarily arises from the sub-problem involving the primal variable, and a first-order dynamical system is simpler and more tractable to solve than a second-order one, He et al. hehufang2022 proposed the following “second-order primal” + “first-order dual” dynamical system for solving Problem (3):
@@
where $\gamma>0$ and $\delta>0$. This system can be viewed as an extension of Polyaks heavy ball with friction system in p1964 . Note that “second-order primal” + “first-order dual” dynamical systems and discrete accelerated algorithms for solving Problem (3) have been studied in 2024zhu ; li2025 ; zhufang2025 ; s2025 .
@@ 1.2.1 Dynamical systems with Hessian-driven damping
It is worth noting that the incorporation of Hessian-driven damping into the dynamical system has been a significant milestone in optimization and mechanics, as Hessian-driven damping can effectively mitigate the oscillations. Building on this foundation, many researchers use dynamical systems with Hessian-driven damping for solving the problem (5). Alvarez et al. a2002 first proposed the second-order dynamical system with constant viscous damping and constant Hessian-driven damping:
@@
where $b=1$. They obtained some new convergence properties of the solution trajectory generated by the system (7). Attouch et al. attouch2016 established the system (7) with $a=\frac{\alpha}{t}$, and obtained the $\mathcal{O}\left(\frac{1}{t^{2}}\right)$ convergence rate of the objective residual for solving the problem (5). To further accelerate the convergence rate, Attouch et al. a2022 proposed the system (7) with $a=\frac{\alpha}{t}$ and $b=\beta(t)$, and obtained the $\mathcal{O}\left(\frac{1}{t^{2}\beta(t)}\right)$ convergence rate for the objective residual. Moreover, for leading to Nesterov-type inertial algorithms via standard implicit or explicit schemes, Alecsa et al. Alecsa2021 proposed the second-order dynamical system with implicit Hessian-driven damping:
@@
They showed that inertial algorithms, such as Nesterov’s algorithm, can be obtained via the natural explicit discretization from this dynamical system, and they also obtained the $\mathcal{O}\left(\frac{1}{t^{2}}\right)$ convergence rate of the objective residual along the trajectory generated by this system. For more results on second order dynamical systems with Hessian-driven damping for solving the problem (5), see bot2019 ; bot2021 ; attouch2023 and the references therein.
However, to the best of our knowledge, it appears that there exist only few papers in the literature devoting to the study of primal-dual dynamical systems with Hessian-driven damping in both continuous and discrete time for solving Problem (3). More precisely, He et al. hetian2025 proposed the following “second-order primal” + “first-order dual” dynamical system:
@@
where $t\geq t_{0}$, $\alpha>1$. They obtained the $\mathcal{O}\left(\frac{1}{t\eta(t)}\right)$ convergence rate for the primal-dual gap, feasibility violation and the objective residual along the trajectory generated by this system, and also obtained corresponding convergence rates for proposed primal-dual algorithm by time discretization of this system. Li et al. lihe2025 proposed a “second-order primal”+“second-order dual” dynamical system with implicit Hessian damping. They also established the fast convergence properties of the proposed dynamical system under suitable conditions. From what was mentioned above, we see that primal-dual dynamical systems with Hessian-driven damping for solving convex optimization problems with linear equality constraints have so far received much less attention compared with others. So, the first aim of this paper is to investigate primal-dual dynamical systems with Hessian-driven damping for solving Problem (3).
@@ 1.2.2 Dynamical systems with closed-loop control
The damping coefficients of a closed-loop dynamical system are directly governed by its evolving trajectory, which classifies the system as autonomous. Unlike non-autonomous systems, autonomous ones are generally preferred in practice because they do not depend on the time variable $t$. As far as we know, there are a few papers in the literature devoted to dynamical systems with closed-loop control for the unconstrained convex optimization problem (5). More precisely, Attouch et al. attouch2022 proposed an inertial dynamical system with closed-loop control:
$\ddot{x}(t)+\gamma(t)\dot{x}(t)+\nabla f(x(t))=0,$ (8)
where $t\geq t_{0}$ and $\gamma(t)=\|\dot{x}(t)\|^{p}$ for several values of $p\in\mathbb{R}$. They obtained the weak convergence of the trajectories to optimal solutions. For achieving convergence rate, Lin and Jordan lin2022 introduced a second-order dynamical system with closed-loop control as follows:
@@
where $c>0$, $\theta\in(0,1)$, and $p\geq 1$. They obtained the $\mathcal{O}\left(\frac{1}{t^{\frac{3p+1}{2}}}\right)$ convergence rate for the objective residual along the trajectory generated by the system (9), and provided two algorithmic frameworks from time discretization of this system. Maier et al. m2023 proposed the system (8) with $\gamma(t):=\sqrt{E(t)}$, where $E(t):=f(x(t))-f(x^{*})+\frac{1}{2}\|\dot{x}(t)\|^{2}$, and obtained the $\mathcal{O}\left(\frac{1}{t^{2-\delta}}\right)$ convergence rate for the objective residual along the trajectory generated by this system with $\delta>0$. By using time scaling and averaging technique, Attouch et al. attouchbot2025 developed autonomous inertial dynamical systems which involve vanishing viscous damping and implicit Hessian-driven damping. They obtained the $\mathcal{O}\left(\frac{1}{t^{1+q-\frac{1}{p}}}\right)$ convergence rate for the objective residual along the trajectory generated by this system with $q>0$, $p\geq 1$ and the weak convergence of the trajectories to optimal solutions.
We observe that there is no results dealing with primal-dual dynamical systems with closed-loop control for solving Problem (3). So, the second aim of this paper is to investigate primal-dual dynamical systems with closed-loop control for solving Problem (3) in continuous and discrete time.
@@ 1.3 Main contributions
The continuous level: We propose a new “second-order primal” + “first-order dual” dynamical system (4) with time scaling and Hessian-driven damping. The novelty of System (4) considered in this paper when compared with the system introduced in hehufang2022 is that System (4) incorporates Hessian-driven damping and all damping coefficients governed by the gradient of the Lagrangian function with a feedback control. We note that this is the first paper for the investigation of primal-dual dynamical systems with closed-loop control for solving convex optimization problems with linear equality constraints. Further, the convergence rate of the objective function residual obtained in this paper matches that reported in lin2022 , which addresses System (4) for unconstrained convex optimization problems.
The discrete level: We develop an accelerated autonomous primal-dual algorithm, where the step size is adaptively defined based on gradient of the objective function for Problem (3). This makes our numerical scheme much easier implementable than the numerical algorithm proposed in lin2022 . By appropriately adjusting these parameters, we show that this algorithm proposed in this paper exhibits the $\mathcal{O}\left(\frac{1}{k^{\frac{3p-1}{2p}}}\right)$ convergence rate for the primal-dual gap, the feasibility violation, and the objective residual. Moreover, in numerical experiments, our algorithm yields a significantly faster convergence rate and consistently achieves substantially higher accuracy than state-of-the-art methods.
@@ 1.4 Organization
The rest of this paper is organized as follows. In Section 2, we obtain fast convergence rates of the primal-dual gap, the feasibility violation, and the objective residual along the trajectories generated by System (4). In Section 3, we propose an accelerated autonomous primal-dual algorithm for solving Problem (3) and give some convergence analysis. In Section 4, we present some numerical experiments to illustrate the obtained results.
@@ 2 The continuous-time dynamical system
In this section, by using the Lyapunov analysis, we establish the fast convergence rates for the primal-dual gap, the feasibility violation, and the objective residual along the trajectory generated by System (4) under mild assumptions on the parameters. To begin, we recall the following result which will play an important role in the sequel.
@@
(attouchbot2025, , Lemma 2.2) Suppose that there exist $C_{0}>0$ and $b>a\geq 0$ such that
@@
where $\tau$ is defined in terms of $\mu$ as in System (4), namely, $\tau(t)=\frac{1}{q^{q}}\left(t_{0}+\int_{t_{0}}^{t}[\mu(s)]^{\frac{1}{q}}ds\right)^{q}$. Then, there exists $C_{1}>0$ such that
@@
$\mathcal{L}(x(t),\lambda^{*})-\mathcal{L}(x^{*},\lambda^{*})=\mathcal{O}\left(\frac{1}{t^{\frac{2qp-p+1}{2}}}\right),~\textup{as}~t\to+\infty,$
$\|Ax(t)-b\|=\mathcal{O}\left(\frac{1}{t^{\frac{2qp-p+1}{2}}}\right),~\textup{as}~t\to+\infty,$
$|f(x(t))-f(x^{*})|=\mathcal{O}\left(\frac{1}{t^{\frac{2qp-p+1}{2}}}\right),~\textup{as}~t\to+\infty,$
@@
Furthermore, $(x(s),\lambda(s))$ is bounded over $t_{0},t]$ and the upper bound only depends on the initial condition.
@@
For any fixed $(x^{*},\lambda^{*})\in\mathcal{S}$, we introduce the energy function $\mathcal{E}:\left[t_{0},+\infty\right)\to\mathbb{R}$ which defined as
@@
This together with (11) follows that
@@
where the third equality holds due to $\nabla_{x}\mathcal{L}\left(x(t),\lambda(t)\right)=\nabla_{x}\mathcal{L}\left(x(t),\lambda^{*}\right)+A^{*}(\lambda(t)-\lambda^{*})$ and the last equality holds due to the third equation of (10) and
@@
From the second equation of the system (10), we have
@@
This implies $\mathcal{E}(t)\leq\mathcal{E}(t_{0})$ for all $t\geq t_{0}$. By the definition of $\mathcal{E}(t)$, we obtain that
$\mathcal{L}\left(x(t),\lambda^{*}\right)-\mathcal{L}\left(x^{*},\lambda^{*}\right)\leq\frac{\mathcal{E}(t_{0})}{\tau(t)}=\mathcal{O}\left(\frac{1}{\tau(t)}\right),~\textup{as}~t\to+\infty,$
and $\lambda(t)$ is bounded on $t_{0},+\infty)$ and the upper bound only depends on the initial condition. From the third equation of the system (10), we have
@@
which means that
$\|Ax(t)-b\|\leq\mathcal{O}\left(\frac{1}{\tau(t)}\right),~\textup{as}~t\to+\infty.$
@@
$|f(x(t))-f(x^{*})|\leq\mathcal{L}(x(t),\lambda^{*})-\mathcal{L}(x^{*},\lambda^{*})+\|\lambda^{*}\|\|Ax(t)-b\|\leq\mathcal{O}\left(\frac{1}{\tau(t)}\right),~\textup{as}~t\to+\infty.$
@@
From the fourth equation of the system (10), we have
@@
This together with the last equation of the system (10) yields
@@
Together with (14), (16) and Lemma 1, there exists $C_{1}>0$ such that
@@
$\mathcal{L}(x(t),\lambda^{*})-\mathcal{L}(x^{*},\lambda^{*})=\mathcal{O}\left(\frac{1}{t^{\frac{2qp-p+1}{2}}}\right),~\textup{as}~t\to+\infty,$
$\|Ax(t)-b\|=\mathcal{O}\left(\frac{1}{t^{\frac{2qp-p+1}{2}}}\right),~\textup{as}~t\to+\infty,$
@@
$|f(x(t))-f(x^{*})|=\mathcal{O}\left(\frac{1}{t^{\frac{2qp-p+1}{2}}}\right),~\textup{as}~t\to+\infty.$
Now, we only need to show that $\tau(s)$ is bounded over $t_{0},t]$ and the upper bound only depends on the initial condition. In fact, by using the definition of $\mathcal{E}(t)$ and $\mathcal{E}(t)\leq\mathcal{E}(t_{0})$, we have
@@
This follows that $v(t)$ is bounded over $t_{0},+\infty)$ and the upper bound only depends on the initial condition. Note that
@@
where the last inequality holds due to the second equation of system (10). From (15) and the last equation of system (10), we have
@@
Combined with (18), (19), $\|v(t)-x^{*}\|\leq\sqrt{2\mathcal{E}(t_{0})}$ and the fact that $\tau(t)$ is monotonically increasing, we have
@@
Now, we analyze (21) in terms of the following two cases.
$\mathbf{Case~I:}$ $1\leq q<2$ and $1\leq p\leq\frac{2}{2-q}$. In this case, from Hölder inequality and the fact that $\tau(t)$ is monotonically increasing, there exists $C_{2}>0$ such that
@@
$\mathbf{Case~II:}$ $q\geq 2$ and $p\geq 1$. In this case, from Hölder inequality and the fact that $\tau(t)$ is monotonically increasing, there exists $C_{3}>0$ such that
@@
From both cases, it follows that $x(s)$ is bounded over $t_{0},t]$ and the upper bound only depends on the initial condition.
@@
Note that Attouch et al. attouchbot2025 introduced a second-order dynamical system via closed-loop control of the gradient for solving the problem (5). In (attouchbot2025, , Theorem 3.5), they obtained the ${o}\left(\frac{\ln(t)}{t^{pq}}\right)$ convergence rate for the objective residual along the trajectory generated by the system. Theorem 2.1 extends (attouchbot2025, , Theorem 3.5) from the unconstrained optimization problem to Problem (5), and achieves a faster convergence rate for the objective residual.
@@
$\mathcal{L}(x(t),\lambda^{*})-\mathcal{L}(x^{*},\lambda^{*})=\mathcal{O}\left(\frac{1}{t^{\frac{p+1}{2}}}\right),~\textup{as}~t\to+\infty,$
$\|Ax(t)-b\|=\mathcal{O}\left(\frac{1}{t^{\frac{p+1}{2}}}\right),~\textup{as}~t\to+\infty,$
$|f(x(t))-f(x^{*})|=\mathcal{O}\left(\frac{1}{t^{\frac{p+1}{2}}}\right),~\textup{as}~t\to+\infty,$
@@
Furthermore, $(x(s),\lambda(s))$ is bounded over $t_{0},t]$ and the upper bound only depends on the initial condition.
@@
$\mathcal{L}(x(t),\lambda^{*})-\mathcal{L}(x^{*},\lambda^{*})=\mathcal{O}\left(\frac{1}{t^{\frac{3p+1}{2}}}\right),~\textup{as}~t\to+\infty,$
$\|Ax(t)-b\|=\mathcal{O}\left(\frac{1}{t^{\frac{3p+1}{2}}}\right),~\textup{as}~t\to+\infty,$
$|f(x(t))-f(x^{*})|=\mathcal{O}\left(\frac{1}{t^{\frac{3p+1}{2}}}\right),~\textup{as}~t\to+\infty,$
@@
Furthermore, $(x(s),\lambda(s))$ is bounded over $t_{0},t]$ and the upper bound only depends on the initial condition.

### v2

# arXiv:2602.16402v2[math.OC] 28 Feb 2026

@@ Abstract
This paper develops a primal-dual dynamical system where the coefficients are designed in closed-loop manner for solving a convex optimization problem with linear equality constraints. We first introduce a “second-order primal” + “first-order dual” continuous-time dynamical system, in which both the time scaling and Hessian-driven damping are governed by a feedback control of the gradient for the Lagrangian function. This system achieves the fast convergence rates for the primal-dual gap, the feasibility violation, and the objective residual along its trajectory. Subsequently, by time discretization of this system, we develop an accelerated primal-dual algorithm with a gradient-defined adaptive step size. We also obtain convergence rates for the primal-dual gap, the feasibility violation, and the objective residual. Furthermore, we provide numerical results to demonstrate the practical efficacy and superior performance of the proposed algorithm.
@@ 1.2 Motivation and related works
Over the past decades, a series of studies have explored second-order dynamical systems in continuous and discrete time for solving the following unconstrained convex optimization problem:
@@
where $\alpha\geq 3$, and obtained an $O\left(\frac{1}{t^{2}}\right)$ convergence rate of the objective residual along the trajectory generated by System (6). System (6) has further been studied in several papers, including may2017 ; attouch2018 ; attouch2019 . They showed that the objective residual converges at a rate of order $o\left(\frac{1}{t^{2}}\right)$ for $\alpha>3$, and the trajectory weakly converges to the global minimizer of the objective function. In order to improve the convergence rate, Attouch et al. a2019 proposed System (6) with time scaling:
@@
where $\alpha\geq 1$ and $\beta:[t_{0},+\infty)\rightarrow(0,+\infty)$ is the time scaling function which is non-decreasing and continuously differentiable. They obtained an $O\left(\frac{1}{t^{2}\beta(t)}\right)$ convergence rate for the objective residual along the trajectory generated by this system, and proposed an inertial proximal algorithm by time discretization of this system.
On the other hand, primal-dual dynamical system methods have also attracted increasing interest among many researchers for solving the linear equality constrained convex optimization problem (3). Over the past few years, there have been a wide variety of works devoted to “second-order primal” + “second-order dual” continuous-time dynamical systems for solving Problem (3), see bot2021n ; hehu2021 ; zeng2023 ; hn2023 . By time discretization of continuous-time dynamical systems, new accelerated primal-dual algorithms for solving Problem (3) have been proposed in hehu2022 ; bot2023 ; zhu2024 ; dingliu2025 . Since the computational cost of a primal-dual method primarily arises from the sub-problem involving the primal variable, and since a first-order dynamical system is simpler and more tractable to solve than a second-order one, He et al. hehufang2022 proposed the following “second-order primal” + “first-order dual” dynamical system for solving Problem (3):
@@
where $\gamma>0$ and $\delta>0$. This system can be viewed as an extension of Polyaks heavy ball with friction system in p1964 . Note that “second-order primal” + “first-order dual” dynamical systems, as well as discrete accelerated algorithms, for solving Problem (3) have been studied in 2024zhu ; li2025 ; zhufang2025 ; s2025 .
@@ 1.2.1 Dynamical systems with Hessian-driven damping
It is worth noting that the incorporation of Hessian-driven damping into the dynamical system has been a significant milestone in optimization and mechanics, as Hessian-driven damping can effectively mitigate the oscillations. Building on this foundation, many researchers use dynamical systems with Hessian-driven damping to solve Problem (5). Alvarez et al. a2002 first proposed the second-order dynamical system with constant viscous damping and constant Hessian-driven damping:
@@
where $b=1$. They obtained some new convergence properties of the solution trajectory generated by System (7). Attouch et al. attouch2016 established System (7) with $a=\frac{\alpha}{t}$, and obtained the $O\left(\frac{1}{t^{2}}\right)$ convergence rate of the objective residual for solving Problem (5). To further accelerate the convergence rate, Attouch et al. a2022 proposed System (7) with $a=\frac{\alpha}{t}$ and $b=\beta(t)$, and obtained the $O\left(\frac{1}{t^{2}\beta(t)}\right)$ convergence rate for the objective residual. Moreover, to derive Nesterov-type inertial algorithms via standard implicit or explicit schemes, Alecsa et al. Alecsa2021 proposed the second-order dynamical system with implicit Hessian-driven damping:
@@
They showed that inertial algorithms, such as Nesterov’s algorithm, can be obtained via the natural explicit discretization of this dynamical system, and they also obtained the $O\left(\frac{1}{t^{2}}\right)$ convergence rate of the objective residual along the trajectory generated by this system. For more results on second order dynamical systems with Hessian-driven damping for solving Problem (5), see bot2019 ; bot2021 ; attouch2023 and the references therein.
However, to the best of our knowledge, it appears that there exist only a few papers in the literature devoted to the study of primal-dual dynamical systems with Hessian-driven damping in both continuous and discrete time for solving Problem (3). More precisely, He et al. hetian2025 proposed the following “second-order primal” + “first-order dual” dynamical system:
@@
where $t\geq t_{0}$, $\alpha>1$. They obtained an $O\left(\frac{1}{t\eta(t)}\right)$ convergence rate for the primal-dual gap, feasibility violation and the objective residual along the trajectory generated by this system, and also obtained the corresponding convergence rates for the proposed primal-dual algorithm by time discretization of this system. Li et al. lihe2025 proposed a “second-order primal”+“second-order dual” dynamical system with implicit Hessian damping. They also established the fast convergence properties of the proposed dynamical system under suitable conditions. From what was mentioned above, we see that primal-dual dynamical systems with Hessian-driven damping for solving convex optimization problems with linear equality constraints have so far received much less attention compared to other types of dynamical systems. Therefore, the first aim of this paper is to investigate such systems for solving Problem (3).
@@ 1.2.2 Dynamical systems with closed-loop control
The damping coefficients of a closed-loop dynamical system are directly governed by its evolving trajectory, which classifies the system as autonomous. Unlike non-autonomous systems, autonomous ones are generally preferred in practice because they do not depend explicitly on the time variable $t$. To the best of our knowledge, only a few papers in the literature have studied dynamical systems with closed-loop control for the unconstrained convex optimization problem (5). More specifically, Lin and Jordan lin2022 introduced a second-order dynamical system with closed-loop control as follows:
@@
where $c>0$, $\theta\in(0,1)$, and $p\geq 1$. They obtained an $O\left(\frac{1}{t^{\frac{3p+1}{2}}}\right)$ convergence rate for the objective residual along the trajectory generated by System (8), and provided two algorithmic frameworks from time discretization of this system. Attouch et al. attouch2022 proposed an inertial dynamical system:
$\ddot{x}(t)+\mathcal{G}\left(\dot{x}(t),\nabla f(x(t)),\nabla^{2}f(x(t))\right)+\nabla f(x(t))=0,$
where $t\geq t_{0}$ and $\mathcal{G}\left(\dot{x}(t),\nabla f(x(t)),\nabla^{2}f(x(t))\right)$ acts as a closed-loop control. They analyzed the asymptotic convergence and the convergence rates of the trajectories generated by this system. Maier et al. m2023 proposed a second-order dynamical system:
$\ddot{x}(t)+\sqrt{E(t)}\dot{x}(t)+\nabla f(x(t))=0,$
where $E(t):=f(x(t))-f(x^{*})+\frac{1}{2}\|\dot{x}(t)\|^{2}$. They obtained an $O\left(\frac{1}{t^{2-\delta}}\right)$ convergence rate for the objective residual along the trajectory generated by this system with $\delta>0$. By using time scaling and averaging technique, Attouch et al. attouchbot2025 developed autonomous inertial dynamical systems which involve vanishing viscous damping and implicit Hessian-driven damping. They obtained an $O\left(\frac{1}{t^{1+q-\frac{1}{p}}}\right)$ convergence rate for the objective residual along the trajectory generated by this system with $q>0$, $p\geq 1$ and the weak convergence of the trajectories to optimal solutions.
We observe that there are no results on primal-dual dynamical systems with closed-loop control for solving Problem (3). Hence, the second aim of this paper is to investigate such systems in both continuous and discrete time.
@@ 1.3 Main contributions
The continuous level: We propose a new “second-order primal” + “first-order dual” dynamical system (4) with time scaling and Hessian-driven damping. The novelty of System (4) considered in this paper, when compared with the system introduced in hehufang2022 , is that System (4) incorporates Hessian-driven damping, with all damping coefficients governed by the gradient of the Lagrangian function with a feedback control. We note that this is the first paper to investigate primal-dual dynamical systems with closed-loop control for solving convex optimization problems with linear equality constraints. Further, the convergence rate of the objective function residual obtained in this paper matches that reported in lin2022 , which addresses System (4) for unconstrained convex optimization problems.
The discrete level: We develop an accelerated autonomous primal-dual algorithm, where the step size is adaptively defined based on gradient of the Lagrangian function. This makes our numerical scheme much easier to implement than the numerical algorithm proposed in lin2022 . By appropriately adjusting the parameters, we show that the proposed algorithm achieves an $O\left(\frac{1}{k^{\frac{3p-1}{2p}}}\right)$ convergence rate for the primal-dual gap, the feasibility violation, and the objective residual. Moreover, in numerical experiments, our algorithm yields a significantly faster convergence rate and consistently achieves substantially higher accuracy than state-of-the-art methods.
@@ 1.4 Organization
The rest of this paper is organized as follows. In Section 2, we establish fast convergence rates for the primal-dual gap, the feasibility violation, and the objective residual along the trajectories generated by System (4). In Section 3, we propose an accelerated autonomous primal-dual algorithm for solving Problem (3) and provide detailed convergence analysis. In Section 4, we present numerical experiments to illustrate the obtained results.
@@ 2 The continuous-time dynamical system
In this section, using the Lyapunov analysis, we establish fast convergence rates for the primal-dual gap, the feasibility violation, and the objective residual along the trajectory generated by System (4) under mild assumptions on the parameters. To begin, we recall the following result, which will play an important role in the sequel.
@@
(attouchbot2025, , Lemma 2.2) Suppose that there exist constants $C_{0}>0$ and $b>a\geq 0$ such that
@@
where $\tau$ is defined in terms of $\mu$ as in System (4), namely, $\tau(t)=\frac{1}{q^{q}}\left(t_{0}+\int_{t_{0}}^{t}[\mu(s)]^{\frac{1}{q}}ds\right)^{q}$. Then, there exists a constant $C_{1}>0$ such that
@@
$\mathcal{L}(x(t),\lambda^{*})-\mathcal{L}(x^{*},\lambda^{*})=O\left(\frac{1}{t^{\frac{2qp-p+1}{2}}}\right),~\textup{as}~t\to+\infty,$
$\|Ax(t)-b\|=O\left(\frac{1}{t^{\frac{2qp-p+1}{2}}}\right),~\textup{as}~t\to+\infty,$
$|f(x(t))-f(x^{*})|=O\left(\frac{1}{t^{\frac{2qp-p+1}{2}}}\right),~\textup{as}~t\to+\infty,$
@@
Furthermore, $(x(s),\lambda(s))$ is bounded over $t_{0},t]$, and the upper bound depends only on the initial condition.
@@
For any fixed $(x^{*},\lambda^{*})\in\mathcal{S}$, we introduce the energy function $\mathcal{E}:\left[t_{0},+\infty\right)\to\mathbb{R}$ defined as follows:
@@
This together with (10) gives
@@
where the third equality holds because $\nabla_{x}\mathcal{L}\left(x(t),\lambda(t)\right)=\nabla_{x}\mathcal{L}\left(x(t),\lambda^{*}\right)+A^{*}(\lambda(t)-\lambda^{*})$, and the last equality follows from the third equation of System (9) and
@@
From the second equation of System (9), we have
@@
This implies $\mathcal{E}(t)\leq\mathcal{E}(t_{0})$ for all $t\geq t_{0}$. By the definition of $\mathcal{E}(t)$, we obtain
$\mathcal{L}\left(x(t),\lambda^{*}\right)-\mathcal{L}\left(x^{*},\lambda^{*}\right)\leq\frac{\mathcal{E}(t_{0})}{\tau(t)}=O\left(\frac{1}{\tau(t)}\right),~\textup{as}~t\to+\infty,$
and $\lambda(t)$ is bounded on $t_{0},+\infty)$, and the upper bound depends only on the initial condition. From the third equation of System (9), it follows that
@@
and hence,
$\|Ax(t)-b\|\leq O\left(\frac{1}{\tau(t)}\right),~\textup{as}~t\to+\infty.$
@@
$|f(x(t))-f(x^{*})|\leq\mathcal{L}(x(t),\lambda^{*})-\mathcal{L}(x^{*},\lambda^{*})+\|\lambda^{*}\|\|Ax(t)-b\|\leq O\left(\frac{1}{\tau(t)}\right),~\textup{as}~t\to+\infty.$
@@
From the fourth equation of System (9), we have
@@
This together with the last equation of System (9) yields
@@
By (13), (15) and Lemma 1, it follows that there exists a constant $C_{1}>0$ such that
@@
$\mathcal{L}(x(t),\lambda^{*})-\mathcal{L}(x^{*},\lambda^{*})=O\left(\frac{1}{t^{\frac{2qp-p+1}{2}}}\right),~\textup{as}~t\to+\infty,$
$\|Ax(t)-b\|=O\left(\frac{1}{t^{\frac{2qp-p+1}{2}}}\right),~\textup{as}~t\to+\infty,$
@@
$|f(x(t))-f(x^{*})|=O\left(\frac{1}{t^{\frac{2qp-p+1}{2}}}\right),~\textup{as}~t\to+\infty.$
Now, we only need to show that $\tau(s)$ is bounded over $t_{0},t]$, and the upper bound depends only on the initial condition. In fact, by using the definition of $\mathcal{E}(t)$ and the fact that $\mathcal{E}(t)\leq\mathcal{E}(t_{0})$, we obtain
@@
Thus, it follows that $v(t)$ is bounded over $t_{0},+\infty)$, and the upper bound depends only on the initial condition. Note that
@@
where the last inequality holds because of the second equation of System (9). From (14) and the last equation of System (9), we have
@@
Thus, by the above inequality, (17), $\|v(t)-x^{*}\|\leq\sqrt{2\mathcal{E}(t_{0})}$ and the fact that $\tau(t)$ is monotonically increasing, it follows that
@@
Now, we analyze (20) by considering the following two cases.
$\mathbf{Case~I:}$ $1\leq q<2$ and $1\leq p\leq\frac{2}{2-q}$. In this case, from Hölder inequality and the fact that $\tau(t)$ is monotonically increasing, it follows that there exists a constant $C_{2}>0$ such that
@@
$\mathbf{Case~II:}$ $q\geq 2$ and $p\geq 1$. In this case, from Hölder inequality and the fact that $\tau(t)$ is monotonically increasing, it is clear that there exists a constant $C_{3}>0$ such that
@@
From both cases, it follows that $x(s)$ is bounded over $t_{0},t]$, and the upper bound depends only on the initial condition.
@@
Note that Attouch et al. attouchbot2025 introduced a second-order dynamical system via closed-loop control of the gradient for solving Problem (5). In (attouchbot2025, , Theorem 3.5), they obtained an ${o}\left(\frac{\ln(t)}{t^{pq}}\right)$ convergence rate for the objective residual along the trajectory generated by the system. Theorem 2.1 extends (attouchbot2025, , Theorem 3.5) from the unconstrained optimization problem to Problem (5), and achieves a faster convergence rate for the objective residual.
@@
$\mathcal{L}(x(t),\lambda^{*})-\mathcal{L}(x^{*},\lambda^{*})=O\left(\frac{1}{t^{\frac{p+1}{2}}}\right),~\textup{as}~t\to+\infty,$
$\|Ax(t)-b\|=O\left(\frac{1}{t^{\frac{p+1}{2}}}\right),~\textup{as}~t\to+\infty,$
$|f(x(t))-f(x^{*})|=O\left(\frac{1}{t^{\frac{p+1}{2}}}\right),~\textup{as}~t\to+\infty,$
@@
Furthermore, $(x(s),\lambda(s))$ is bounded over $t_{0},t]$, and the upper bound depends only on the initial condition.
@@
$\mathcal{L}(x(t),\lambda^{*})-\mathcal{L}(x^{*},\lambda^{*})=O\left(\frac{1}{t^{\frac{3p+1}{2}}}\right),~\textup{as}~t\to+\infty,$
$\|Ax(t)-b\|=O\left(\frac{1}{t^{\frac{3p+1}{2}}}\right),~\textup{as}~t\to+\infty,$
$|f(x(t))-f(x^{*})|=O\left(\frac{1}{t^{\frac{3p+1}{2}}}\right),~\textup{as}~t\to+\infty,$
@@
Furthermore, $(x(s),\lambda(s))$ is bounded over $t_{0},t]$, and the upper bound depends only on the initial condition.